please help me to solve this question.
The base of a triangular field is three times its altitude. If the cost of sowing the field at Rs 58 per hectare is Rs 783, find its base and height.

Answer by LinnW(1048)

783/58 = 13.5 number of hectares
one hectare is 10,000 square meters so
783/58 = 13.5*10000 square meters = 135000 square meters
The area of the triangle is 0.5*base*height
Set b = base
a = altitude
b = 3a
So 0.5*b*a = 135000
Substitute 3a for b
0.5*3a*a = 135000
1.5a^2 = 135000
divide each side by 1.5
a^2 = 135000/1.5
a^2 = 90000
a = 300 meters
b = 900 meters